OpenStudy (anonymous):

HELP!!!!! AND SHOW WORK Solve the following system. y = x + 3 3x + y = 19 (7, 4) (−4, 7) (4, 7) (4, −7)

OpenStudy (anonymous):

y= x+3 you can plug the (y) in the other equation you will get 3x +x+3 = 19 4x+3= 19 4x =19-3 4x = 16 x = 4 you now can get y by puging in x to the equation y = 4+3 y= 7

OpenStudy (anonymous):

Thea answers are (7, 4) (−4, 7) (4, 7) (4, −7)